A Chemist ahs 25% and a 50% acid soultion. How much of each soultion should be used to form 200 ml of a 35% acid soultion.

Answer by checkley71(8403)

.25X+.50(200-X)=200*.35
.25X+100-.50X=70
-.25X=70-100
-.25X=-30
X=-30/-.25
X=120 ML OF 25% ACID IS REQUIRED.
200-120=80 ML OF 50% ACID IS ALSO REQUIRED.
PROOF
.25*120+.50*80=70
30+40=70
70=70